OpenStudy (sushi121212):

A grocer buys cases of almonds and walnuts. Almonds are packaged 20 bags per case. The grocer pays $30 per case of almonds and makes a profit of $17 per case. Walnuts are packaged 24 bags per case.The grocer pays $26 per case of walnuts and makes a profit of $15 per case. He orders no more than 300 bags of almonds and walnuts together at a maximum "cost" of $400.

OpenStudy (sushi121212):

a. Write the constraints. Use X for the # of cases of almonds ordered and Y for the number of cases of walnut b. Graph the constraints c. Write the objective function for the profit d. How many cases of almonds and walnuts maximize the grocer's profit?

OpenStudy (sushi121212):

I'm very confused on how to write my constraints.

OpenStudy (phi):

you start by using Use X for the # of cases of almonds ordered and Y for the number of cases of walnut x is the number of cases of almonds each case of almonds has 20 bags. Now consider: if you had 1 case of almonds, you would have 20 bags if you had 2 cases, then 40 bags or 3 cases, then 60 bags notice we do: number of cases times 20 use that rule, but use x as the "number" of cases can you write how many bags are in x cases of almonds?

OpenStudy (sushi121212):

Is it 20x?

OpenStudy (phi):

yes. and there are 24y bags in y cases of almonds

OpenStudy (phi):

if we add up all the bags, we need no more than 300 bags any idea how to write that ?

OpenStudy (sushi121212):

x+y≤300?

OpenStudy (phi):

almost. but x is the number of *cases* of almonds, and y is cases of walnuts we want *bags* (which we figured out up above)

OpenStudy (sushi121212):

Oh! So it's 20x+24y≤300?

OpenStudy (phi):

yes and if we write it in "slope-intercept form" y <= (-20/24)y +300/24 y <= (-5/6)y + 12.5 which we can plot

OpenStudy (sushi121212):

I'm a bit confused... Why is there 2 y's for each equation?

OpenStudy (phi):

oh, bad typing. it should be y <= (-5/6)x + 12.5

OpenStudy (sushi121212):

Do I plug in my answers after that? Or is there another equation?

OpenStudy (phi):

this is a complicated question (lots of steps). let's do the cost. pays $30 per case of almonds (x) pays $26 per case of walnuts (y) and idea how to write the total cost for x cases of almonds and y cases of walnuts ?

OpenStudy (phi):

it is almost the same idea as for the number of bags

OpenStudy (sushi121212):

30x+26y≤400?

OpenStudy (phi):

yes, do you have a way to plot these ?

OpenStudy (sushi121212):

By using a graph?

OpenStudy (phi):

we did part a 20x+24y≤300 30x+26y≤400 now part b says plot them. So you have to use either graph paper or a computer

OpenStudy (sushi121212):

There is a piece of graph paper on the worksheet. It's numbered up until 20 on both the x and y axis, but I don't have a picture of it.

OpenStudy (phi):

ok, it should look like this

OpenStudy (sushi121212):

Oh! Okay, thank you! So what do I do after this?

OpenStudy (phi):

now write down the profit same idea as the constraints

OpenStudy (sushi121212):

Does that mean the points on the graph?

OpenStudy (phi):

makes a profit of $17 per case of almonds (x) $15 per case of walnuts (y)

OpenStudy (sushi121212):

17x+15y=P?

OpenStudy (phi):

yes now to find the best profit, you test the points on the graph (the "corners") for example, test (0,12.5). The profit is 17*0 +15*12.5= 187.50 test the other points, and pick the one with the best profit.

OpenStudy (sushi121212):

(9, 5) 17*9+15*5=228 (13.33, 0) 17*13.33+ 15*0=226.61 Is (9, 5) the answer?

OpenStudy (phi):

looks like it so 9 cases of x (almonds) and 5 cases of y (walnuts)

OpenStudy (sushi121212):

So the maximum profit is 9 cases of almonds and 5 cases of walnuts? And that's the end?

OpenStudy (sushi121212):

Thank you so much!! I wish I could give you 100,000,000 medals for helping me for so long!